Evaluate the function.
Given that f(x) = 3 - 5x, find f(-r).
A) 3 + 5r
B) 3 + rx
C) 3 - 5r
D) r - 5x

Answer by Alan3354(69443)

Given that f(x) = 3 - 5x, find f(-r)
Sub -r for x
f(-r) = 3 - 5*(-r)
f(-r) = 3 + 5r